Archimedes Principle Pre-Lab Completely READ the earchimedes principal A cylindrical body of diameter d heigh h and mass m is completely immersed in fluid of density pf
Answer #1

1) Density of the body is given by :

4m 7m

2) Bouyant force acting on m is given by :

gpfT

3) Apparent weight of the body = Actual weight of the body - Weight of liquid displaced

\therefore W_f = mg - \rho_f V_f g = mg - \frac{\rho_f \pi d^2 hg}{4}

4) Bouyancy or bouyant force is given by :

F= (\rho_f-\rho) V_fg =\rho_fV_fg - mg =-W_f

5)

Density of body Density of fluid Weight of body Weight of fluid

or, \frac{\rho}{\rho_f} = \frac{mg}{mg - W_f} \\\Rightarrow \boxed{\rho = \frac{mg\rho_f}{mg - W_f}}